OpenStudy (anonymous):

Melinda has nickels and quarters in her bank. She has four fewer nickels than quarters. She has $3.70 in the bank. How many coins of each does she have?

OpenStudy (anonymous):

To solve this you need to set up a system of equations. In this case these would be your two equations: n+4=q .25q+.05n=3.70 Now you should be able to solve from here!

OpenStudy (shamuta):

I thought it would be 4-Q=N since there are four fewer nickels than quarters

OpenStudy (shamuta):

^make sure to replace the numbers with the correct cents.

OpenStudy (anonymous):

Shamuta, I don't think that's correct. She has four fewer nickels than quarters. Therefore she has four more quarters than nickels so it's either q-4=n or n+4=q